Sarita signed up with Netflix for a fixed fee of $16.99 per month. For this fee, she can receive up to 3
DVDs at a time in the mail and exchange each DVD as often as she likes. She also receives unlimited
instant access to movies being streamed from Netflix to her computer or TV. During the average month in
2010, Sarita received and watched 6 movies sent to her through the mail and she watched an additional 13
movies which were streamed to her computer. What is the average cost of a movie to Sarita? What is the
marginal cost of an additional movie?
1
Expert's answer
2020-05-03T17:10:54-0400

Solution


Given:

Fixed fee per month = $16.99

quantity of DVD received = 3 DVD

additional quantity is 3 DVD


Average cost = "\frac{total fixed fee}{quantity of DVD}"

=total fixed fee per month /quantity of DVD received

=    "\frac{\$16.99}{3}"              

 = $16.99 / 3

= $5.663

This means that the average cost for 1 DVD is $5.663


In 2010, Sarita received and watched 6 movies sent to her through the mail, therefore she need to pay an additional cost for 3 DVD.

It is being said that She also receives unlimited instant access to movies being streamed from Netflix to her computer or TV, for this reason, no additional cost will be charge for 13 movies which were streamed to her computer.  

 

The cost for 6 DVD = "{6}* {\$5.663}" = $33.98

Marginal cost = change in cost/ change in quantity demanded

"\frac{change in cost}{change in quantity demanded}"

                  = "\frac{[\$33.98 - \$16.99]}{6−3}" = "\frac{16.99}{3}" = $5.663